# CKA-Guided Modular Quantization: Beyond Bit-Width to Algorithmic Diversity

### v1

@@ Abstract
Current mainstream post-training quantization methods for large language models typically apply a uniform quantization strategy across all network layers, overlooking the substantial differences in algorithmic suitability among layers. To address this limitation, we propose CKA-Guided Modular Quantization—a fine-tuning-free, plug-and-play framework for algorithmic heterogeneous quantization. Our method independently evaluates multiple PTQ algorithms on each layer and employs Linear Centered Kernel Alignment (CKA) as a metric to automatically select the optimal quantization strategy per layer. The individually optimized strategies are then integrated to construct a hybrid quantized model. Experiments demonstrate that our approach consistently outperforms both uniform quantization baselines and state-of-the-art mixed-precision methods across mainstream LLMs—including LLaMA and Qwen—in terms of perplexity (PPL) and downstream task performance.
CKA-Guided Modular Quantization: Beyond Bit-Width to Algorithmic Diversity
Jinhao Zhang1   Yunquan Zhang2   Daning Cheng2 1Beijing University of Posts and Telecommunications, Beijing, China 2Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China
@@ 1 Introduction
Post-training quantization (PTQ) is a key technique for compressing large language models (LLMs) to enable efficient deployment. Existing PTQ methods—such as GPTQ, AWQ, and SmoothQuant—typically apply the same quantization strategy uniformly across all layers. While this simplifies implementation, it overlooks an important empirical fact: different decoder layers vary significantly in their sensitivity to quantization, and thus benefit differently from distinct quantization approaches.
@@
To implement this idea, we propose CKA-Guided Modular Quantization. We use Linear Centered Kernel Alignment (CKA)—a well-established metric for measuring similarity between neural network representations—as a proxy for functional fidelity. Specifically, for each decoder layer, we apply multiple PTQ methods independently and compute the CKA similarity between the activations of each quantized variant and those of the original full-precision model, using a small calibration dataset. The method yielding the highest CKA score is selected for that layer, and the chosen layers are assembled into a single heterogeneous quantized model.
We evaluate our method on several LLMs, including LLaMA and Qwen, across language modeling and downstream tasks. Results show that our approach consistently outperforms both individual PTQ baselines and state-of-the-art mixed-precision methods—without any retraining or fine-tuning.
@@
  • Extensive experiments show that our method achieves state-of-the-art performance across multiple models and benchmarks.
@@ 2 Related work
Post-Training Quantization Foundational works such as AdaRound (DBLP:conf/icml/NagelABLB20), GPTQ (GPTQ), AWQ (AWQ), and SmoothQuant (SmoothQuant) have established widely adopted paradigms for minimizing quantization error through weight reconstruction and outlier mitigation. Building on these baselines, recent advanced methods introduce novel perspectives: QuIP (DBLP:conf/nips/CheeCKS23) utilizes incoherence processing to enhance extreme low-bit quantization, while OmniQuant (DBLP:conf/iclr/ShaoC0XZLZ00024) introduces learnable clipping thresholds. Furthermore, SpinQuant (spinquant) optimizes rotation matrices to suppress outliers before quantization.
Centered Kernel Alignment Centered Kernel Alignment (CKA)(DBLP:conf/nips/MorcosRB18), particularly its linear variant (CKA), has emerged as a standard proxy for gauging neural representational similarity. It has been widely employed to analyze training dynamics (DBLP:conf/iclr/NguyenRK21), compare diverse architectures (DBLP:conf/nips/RaghuUKZD21), and guide model compression DBLP:conf/iccv/TungM19; DBLP:conf/nips/MaFW23. Notably, high CKA similarity is often interpreted as a signal of redundancy, utilized for expert pruning in Mixture-of-Experts models (MOE_QUANT) and model merging in KnOTS (modelmerge).
@@ 3 Methodology
Figure 2: Overview of our CKA-Guided Modular Quantization Framework. (a) We first analyze layer-wise sensitivity using CKA. (b) Then, we competitively select the optimal quantization method (e.g., GPTQ & SmoothQuant) for each layer. (c) Finally, we integrate these layers into a unified model. This framework achieves optimal heterogeneity at the algorithmic level without retraining.
@@ 3.1 Algorithmic Suitability
Different PTQ algorithms exhibit fundamental differences in their design principles and optimization objectives, leading to varying sensitivities to weight distributions. GPTQ employs a layer-wise greedy optimization approach, formulating quantization error minimization as a constrained quadratic programming problem and updating weights sequentially using Hessian information; it is particularly effective for layers with concentrated weight distributions and few outliers. In contrast, AWQ identifies “sensitive” weight channels—those most influential to the output—based on activation magnitudes and preserves higher precision for them while aggressively quantizing the rest. This method assumes a strong correlation between weight importance and the magnitude of corresponding activations, enabling superior performance when activation distributions are skewed and contain prominent hotspot channels. SmoothQuant, on the other hand, introduces a smoothing factor between weights and activations to shift quantization difficulty from the activation side to the weight side. By reparameterizing the forward pass to balance dynamic ranges, it demonstrates greater robustness in scenarios involving extreme weight outliers and large variations in activation magnitudes.
@@
3.2 CKA-Guided Modular Quantization Framework
Motivated by the observed layer-wise heterogeneity in algorithmic suitability, we propose the CKA-Guided Modular Quantization Framework, whose overall pipeline is illustrated in Algorithm 1 and Figure 2. The core objective of this framework is to allow each layer to select the quantization method that best preserves its representational structure, rather than enforcing a single homogeneous strategy across the entire model. The proposed framework consists of three main stages.
Algorithm 1 CKA-Guided Modular Quantization Framework
@@ 3.2 CKA-Guided Modular Quantization Framework
First, we construct a candidate pool composed of multiple Post-Training Quantization (PTQ) algorithms, such as GPTQ, AWQ, and SmoothQuant. This pool introduces algorithmic diversity, enabling the framework to flexibly handle heterogeneous weight distributions and activation outliers across different layers.
@@ 3.3 Greedy Layer-Wise Selection Strategy
Finding the globally optimal assignment of quantization methods across all layers requires solving a combinatorial optimization problem with a search space of $|\mathcal{P}|^{L}$, which is computationally infeasible for modern large language models. To address this challenge, we adopt a greedy layer-wise selection strategy that optimizes each layer locally, conditioned on the outputs of the preceding quantized layers.
@@ 3.4 Complexity and Inference Efficiency
Offline Search Cost.
The proposed greedy search strategy exhibits linear complexity $O(L\times|\mathcal{P}|)$. For a typical 70B model with $L=80$ layers and a candidate pool of four quantization methods, only $80\times 4=320$ layer-wise evaluations are required. This constitutes a one-time offline cost that is negligible compared to the expense of model pretraining or fine-tuning.
Online Inference Latency.
The proposed framework introduces no additional latency during online inference. Since all layers are constrained to the same bit-width and group size, the memory access pattern remains unchanged. The only theoretical overhead stems from switching between different quantization kernels, which incurs a GPU dispatch cost on the order of microseconds and is therefore negligible compared to the millisecond-level execution time of matrix multiplications.
@@ 4.1 Experiments configuration
Model We evaluate the proposed method on two representative model series: Llama-3-8B, Llama-3.2-1B, Llama-3.2-3B(touvron2023llama), Qwen2.5-1.5B, and Qwen2.5-0.5B(bai2023qwen).
@@
BenchMark We benchmark our approach against SOTA methods, including GPTQ (GPTQ), AWQ (AWQ), SmoothQuant (SmoothQuant), and SpinQuant (spinquant).
@@ 4.2 CKA-Guided Modular Quantization Results
Figure 4: Layer-wise method selection results derived from our CKA-guided framework.
Figure 4 illustrates the distribution of optimal quantization methods automatically selected by our CKA-guided framework for each layer across large language models of varying scales, including Llama8B,Llama3B, Llama1B, Qwen-0.5B, and Qwen-1.5B. It is evident from the figure that, in every model, no single quantization method dominates across the entire network depth. Instead, SmoothQuant (yellow), AWQ (blue), and GPTQ (orange) exhibit a highly interleaved and dynamically shifting allocation pattern.
@@ 4.3 Perplexity Evaluation on LLMs
Ours 12.72 6.89 22.88 11.70 17.41 9.63 16.59 12.87 24.87 14.87
Table 1: Perplexity (PPL) evaluation. We compare our method with FP16 baseline and SOTA PTQ methods.
We evaluated the perplexity (the lower the better) of different quantization methods on multiple LLMs. The methods examined include FP16, AWQ, GPTQ, SmoothQuant, SpinQuant, and our proposed approach (OurS). We utilize the C4 and WikiText-2(Wiki2) to evaluate performance on web text and high-quality articles, respectively.
@@
Model size exhibits a significant influence on quantization robustness. For larger models (e.g., Llama-3-8B), the PPL increase on both C4 and WikiText-2 is relatively modest across all methods, including OurS, indicating that larger models are more tolerant to quantization-induced errors. In contrast, smaller models (e.g., Llama-3.2-1B and Qwen1.5-0.5B) show a more pronounced rise in PPL, suggesting greater sensitivity to quantization.
Our proposed method (OurS) achieves the best or nearly the best PPL performance across all experimental settings. Specifically:
On Llama-3-8B, OurS attains a C4 PPL of 12.72, notably lower than other quantization methods and closest to the FP16 baseline (12.28).
On Qwen1.5-1.5B, OurS reaches a WikiText-2 PPL of 12.87, which is 0.44 lower than the next best quantization method, SpinQuant (13.31).
Even on the very lightweight Qwen1.5-0.5B, OurS maintains a relatively low PPL increase (24.87 on C4), substantially outperforming AWQ (25.98) and GPTQ (26.04).
The experimental results demonstrate that OurS achieves method heterogeneity by adaptively and collaboratively selecting different quantization methods across layers. This optimal method selection, grounded in representational similarity, enables the model to surpass the performance ceiling of any single quantization method. This improvement translates directly into significant performance advantages on language modeling tasks, yielding lower perplexity degradation compared to baseline methods.
@@ 4.4 Downstream experiments
To comprehensively evaluate the capabilities of quantized models in real-world tasks, we conduct a systematic assessment across multiple representative downstream tasks, covering mathematical reasoning (GSM8K), code generation (HumanEval), commonsense reasoning (HellaSwag), and multi-task understanding (MMLU). Each task follows widely accepted evaluation protocols: MMLU (5-shot), GSM8K (8-shot), HellaSwag (10-shot), and HumanEval (0-shot), with accuracy serving as the primary evaluation proxy (higher is better). The experiment compares various quantization methods, including our proposed approach (Ours), to comprehensively reflect the ability of each quantization method to maintain model performance across diverse tasks.
@@
Table  2 presents a quantitative performance comparison of different models on multiple downstream tasks. Generally, all quantization methods result in accuracy degradation compared to the FP16 baseline; however, the magnitude of this decline varies significantly depending on the task type, model scale, and quantization methods. Notably, our proposed method (Ours) maintains optimal or near-optimal performance in most scenarios.
Distinct tasks exhibit varying sensitivities to quantization errors. Mathematical reasoning (GSM8K) and code generation (HumanEval) are the most fragile, showing the steepest accuracy drops across all models. For instance, on Llama-3-8B, accuracy reduction on GSM8K ranges from 2.84 (Ours) to 5.28 (GPTQ), while on HumanEval, the drop ranges from 1.04 (Ours) to 2.26 (GPTQ). In contrast, commonsense reasoning (HellaSwag) and multi-task understanding (MMLU) exhibit relatively minor degradation.
The efficacy of our approach extends to smaller, more sensitive models. On Qwen1.5-0.5B, our method achieves a GSM8K score of 25.12, marking a substantial improvement of 3.25 and 4.90 over AWQ and GPTQ, respectively.
These empirical results strongly validate the effectiveness of the layer-wise adaptive hybrid quantization methods employed by Ours. Our framework effectively identifies and accommodates the heterogeneous quantization method requirements across different model layers, thereby achieving efficient, low-loss quantization while strictly preserving the model’s core inferential capability.
@@ 4.5 Configuration Experiments
To rigorously evaluate the individual contributions of each component and validate the design rationale underlying our CKA-guided modular quantization framework, we conducted extensive ablation studies.
@@ 4.5.1 Impact of PTQ Candidate Diversity
A central motivation of our approach is the premise that a single quantization method cannot optimally accommodate the distinct characteristics of every layer. To validate this, we systematically excluded individual quantization algorithms from the candidate pool and observed the resulting impact on model performance. Table 3 presents our experimental findings.
@@
Table 3: Downstream task performance across five LLMs. Higher accuracy($\uparrow$) indicates better performance.(GSM=GSM8K, HE=HumanEval, HS=HellaSwag).
Experimental results demonstrate that retaining the full spectrum of candidate quantization methods yields superior performance across all evaluated models and tasks. The exclusion of any single algorithm invariably leads to a marked degradation in model fidelity. For instance, on the Llama-3-8B model, our comprehensive strategy pool achieves a C4 PPL of 12.72. In contrast, removing SmoothQuant causes the C4 PPL to deteriorate significantly to 13.55, while excluding AWQ results in a drop in GSM8K reasoning accuracy from 74.33% to 72.85%.

### v2

# CALM: A CKA-Guided Adaptive Layer-Wise Modularization Framework for LLM Quantization

@@ Abstract
Current mainstream post-training quantization methods for large language models typically apply a uniform quantization strategy across all network layers, overlooking the substantial differences in algorithmic suitability among layers. To address this limitation, we propose CALM (A CKA-guided Adaptive Layer-wise Modularization)—a fine-tuning-free, plug-and-play framework for algorithmic heterogeneous quantization. CALM independently evaluates multiple PTQ algorithms on each layer and employs Linear Centered Kernel Alignment (CKA) as a metric to automatically select the optimal quantization strategy per layer. The individually optimized strategies are then integrated to construct a hybrid quantized model. Experiments demonstrate that our approach consistently outperforms both uniform quantization baselines and state-of-the-art mixed-precision methods across mainstream LLMs—including LLaMA and Qwen—in terms of perplexity (PPL) and downstream task performance.
CALM: A CKA-Guided Adaptive Layer-Wise Modularization Framework for LLM Quantization
Jinhao Zhang, Yunquan Zhang, Daning Cheng, Jun Sun, Zicheng Yan Beijing University of Posts and Telecommunications, Beijing, China Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China Zhejiang University, Hangzhou, China University of Science and Technology of China, Hefei, China
@@ 1 Introduction
Post-Training Quantization (PTQ) has emerged as a cornerstone for compressing and deploying Large Language Models (LLMs). Existing PTQ methods—such as GPTQ, AWQ, and SmoothQuant—typically enforce a uniform quantization strategy across all network layers. While this simplifies implementation, it overlooks an important empirical fact: distinct transformer layers exhibit vastly different sensitivities to quantization errors and outlier distributions.
@@
To implement this idea, we propose CKA-guided Adaptive Layer-wise Modularization (CALM). We use Linear Centered Kernel Alignment (CKA)—a well-established metric for measuring similarity between neural network representations—as a proxy for functional fidelity. Specifically, for each decoder layer, we apply multiple PTQ methods independently and compute the CKA similarity between the activations of each quantized variant and those of the original full-precision model, using a small calibration dataset. The method yielding the highest CKA score is selected for that layer, and the chosen layers are assembled into a single heterogeneous quantized model.
We evaluate CALM on several LLMs, including LLaMA and Qwen, across language modeling and downstream tasks. Results show that CALM consistently outperforms both individual PTQ baselines and state-of-the-art mixed-precision methods—without any retraining or fine-tuning.
@@
  • Extensive experiments show that CALM achieves state-of-the-art performance across multiple models and benchmarks.
@@ 2 Related work
Post-Training Quantization Foundational works such as AdaRound (Nagel et al., 2020), GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024), and SmoothQuant (Xiao et al., 2023) have established widely adopted paradigms for minimizing quantization error through weight reconstruction and outlier mitigation. Building on these baselines, recent advanced methods introduce novel perspectives: QuIP (Chee et al., 2023) utilizes incoherence processing to enhance extreme low-bit quantization, while OmniQuant (Shao et al., 2024) introduces learnable clipping thresholds. Furthermore, SpinQuant (Liu et al., 2025) optimizes rotation matrices to suppress outliers before quantization.
Centered Kernel Alignment Centered Kernel Alignment (CKA), particularly its linear variant (Kornblith et al., 2019; Morcos et al., 2018), has emerged as a standard proxy for gauging neural representational similarity. It has been widely employed to analyze training dynamics (Nguyen et al., 2021), compare diverse architectures (Raghu et al., 2021), and guide model compression Tung and Mori (2019); Ma et al. (2023). Notably, high CKA similarity is often interpreted as a signal of redundancy, utilized for expert pruning in Mixture-of-Experts models (Lu et al., 2024) and model merging in KnOTS (Stoica et al., 2025).
@@ 3 Methodology
Figure 2: Overview of the CALM framework (a) We first analyze layer-wise sensitivity using CKA. (b) Then, we competitively select the optimal quantization method (e.g., GPTQ & SmoothQuant) for each layer. (c) Finally, we integrate these layers into a unified model. This framework achieves optimal heterogeneity at the algorithmic level without retraining.
@@ 3.1 Algorithmic Suitability
State-of-the-art PTQ algorithms are built upon distinct and complementary design principles, leading to varying suitability across layers. GPTQ employs a layer-wise greedy optimization approach, formulating quantization error minimization as a constrained quadratic programming problem and updating weights sequentially using Hessian information; it is particularly effective for layers with concentrated weight distributions and few outliers. In contrast, AWQ identifies “sensitive” weight channels—those most influential to the output—based on activation magnitudes and preserves higher precision for them while aggressively quantizing the rest. This method assumes a strong correlation between weight importance and the magnitude of corresponding activations, enabling superior performance when activation distributions are skewed and contain prominent hotspot channels. SmoothQuant, on the other hand, introduces a smoothing factor between weights and activations to shift quantization difficulty from the activation side to the weight side. By reparameterizing the forward pass to balance dynamic ranges, it demonstrates greater robustness in scenarios involving extreme weight outliers and large variations in activation magnitudes.
@@
3.2 The CALM Framework
Motivated by the observed layer-wise heterogeneity in algorithmic suitability, we propose CALM, whose overall pipeline is illustrated in Algorithm 1 and Figure 2. The core objective of this framework is to allow each layer to select the quantization method that best preserves its representational structure, rather than enforcing a single homogeneous strategy across the entire model. The proposed framework consists of three main stages.
Algorithm 1 CALM
@@ 3.2 The CALM Framework
First, we construct a candidate pool composed of multiple Post-Training Quantization (PTQ) algorithms, such as GPTQ, AWQ, and SmoothQuant. While SpinQuant shows high accuracy, we exclude it from the candidate pool to avoid its rotation-based inference overhead, ensuring zero additional latency.
@@ 3.3 Greedy Layer-Wise Selection Strategy
Finding the globally optimal assignment of quantization methods across all layers requires solving a combinatorial optimization problem with a search space of $|\mathcal{P}|^{L}$, which is computationally infeasible for modern large language models. To address this challenge, we adopt a greedy layer-wise selection strategy that optimizes each layer locally. While this greedy approach optimizes locally, extensive experiments confirm it is sufficient to transcend the performance ceilings of monolithic methods. CALM consistently outperforms state-of-the-art (SOTA) uniform quantization baselines across various LLMs, including LLaMA and Qwen."
@@ 3.4 Complexity and Inference Efficiency
Offline Search Cost. The proposed greedy search strategy exhibits linear complexity $O(L\times|\mathcal{P}|)$. For a typical 70B model with $L=80$ layers and a candidate pool of four quantization methods, only $80\times 4=320$ layer-wise evaluations are required. This constitutes a one-time offline cost that is negligible compared to the expense of model pretraining or fine-tuning.
Zero Overhead CALM introduces no additional latency. By finalizing layers as standard W4A8 modules with fixed group sizes, we eliminate dynamic kernel switching. Benchmarks on NVIDIA A800 GPUs confirm negligible latency deviation ( 0.60%, see Appendix A.5), matching the throughput of uniform quantization.
@@ 4.1 Experiments configuration
Model We evaluate the proposed method on two representative model series: Llama-3-8B, Llama-3.2-1B, Llama-3.2-3B(Touvron et al., 2023), Qwen2.5-1.5B, and Qwen2.5-0.5B(Yang et al., 2024).
@@
BenchMark We benchmark CALM against SOTA methods, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024), SmoothQuant (Xiao et al., 2023), and SpinQuant (Liu et al., 2025).
@@ 4.2 CKA-Guided Modular Quantization Results
Figure 4: Layer-wise method selection results derived from the CALM framework.
Figure 4 illustrates the distribution of optimal quantization methods automatically selected by the CALM framework for each layer across large language models of varying scales, including Llama8B,Llama3B, Llama1B, Qwen-0.5B, and Qwen-1.5B. It is evident from the figure that, in every model, no single quantization method dominates across the entire network depth. Instead, SmoothQuant (yellow), AWQ (blue), and GPTQ (orange) exhibit a highly interleaved and dynamically shifting allocation pattern.
@@ 4.3 Perplexity Evaluation on LLMs
CALM 12.72 6.89 22.88 11.70 17.41 9.63 16.59 12.87 24.87 14.87
Table 1: Perplexity (PPL) evaluation. We compare CALM with FP16 baseline and SOTA PTQ methods.
We evaluated the perplexity (the lower the better) of different quantization methods on multiple LLMs. The methods examined include FP16, AWQ, GPTQ, SmoothQuant, SpinQuant, and our proposed approach (CALM). We utilize the C4 and WikiText-2(Wiki2) to evaluate performance on web text and high-quality articles, respectively.
@@
Model size exhibits a significant influence on quantization robustness. For larger models (e.g., Llama-3-8B), the PPL increase on both C4 and WikiText-2 is relatively modest across all methods, including CALM, indicating that larger models are more tolerant to quantization-induced errors. In contrast, smaller models (e.g., Llama-3.2-1B and Qwen1.5-0.5B) show a more pronounced rise in PPL, suggesting greater sensitivity to quantization.
CALM achieves the best or nearly the best PPL performance across all experimental settings. Specifically:
On Llama-3-8B, CALM attains a C4 PPL of 12.72, notably lower than other quantization methods and closest to the FP16 baseline (12.28).
On Qwen1.5-1.5B, CALM reaches a WikiText-2 PPL of 12.87, which is 0.44 lower than the next best quantization method, SpinQuant (13.31).
Even on the very lightweight Qwen1.5-0.5B, CALM maintains a relatively low PPL increase (24.87 on C4), substantially outperforming AWQ (25.98) and GPTQ (26.04).
The experimental results demonstrate that CALM achieves method heterogeneity by adaptively and collaboratively selecting different quantization methods across layers. This optimal method selection, grounded in representational similarity, enables the model to surpass the performance ceiling of any single quantization method. This improvement translates directly into significant performance advantages on language modeling tasks, yielding lower perplexity degradation compared to baseline methods.
@@ 4.4 Downstream experiments
To comprehensively evaluate the capabilities of quantized models in real-world tasks, we conduct a systematic assessment across multiple representative downstream tasks, covering mathematical reasoning (GSM8K), code generation (HumanEval), commonsense reasoning (HellaSwag), and multi-task understanding (MMLU). Each task follows widely accepted evaluation protocols: MMLU (5-shot), GSM8K (8-shot), HellaSwag (10-shot), and HumanEval (0-shot), with accuracy serving as the primary evaluation proxy (higher is better). The experiment compares various quantization methods, including CALM, to comprehensively reflect the ability of each quantization method to maintain model performance across diverse tasks.
Table  2 presents a quantitative performance comparison of different models on multiple downstream tasks. Generally, all quantization methods result in accuracy degradation compared to the FP16 baseline; however, the magnitude of this decline varies significantly depending on the task type, model scale, and quantization methods. Notably, our proposed method (CALM) maintains optimal or near-optimal performance in most scenarios.
Distinct tasks exhibit varying sensitivities to quantization errors. Mathematical reasoning (GSM8K) and code generation (HumanEval) are the most fragile, showing the steepest accuracy drops across all models. For instance, on Llama-3-8B, accuracy reduction on GSM8K ranges from 2.84 (CALM) to 5.28 (GPTQ), while on HumanEval, the drop ranges from 1.04 (CALM) to 2.26 (GPTQ). In contrast, commonsense reasoning (HellaSwag) and multi-task understanding (MMLU) exhibit relatively minor degradation.
The efficacy of CALM extends to smaller, more sensitive models. On Qwen1.5-0.5B, CALM achieves a GSM8K score of 25.12, marking a substantial improvement of 3.25 and 4.90 over AWQ and GPTQ, respectively.
These empirical results strongly validate the effectiveness of the layer-wise adaptive hybrid quantization methods employed by CALM. The CALM framework effectively identifies and accommodates the heterogeneous quantization method requirements across different model layers, thereby achieving efficient, low-loss quantization while strictly preserving the model’s core inferential capability.
@@ 4.5 Configuration Experiments
To rigorously evaluate the individual contributions of each component and validate the design rationale underlying the CALM framework, we conducted extensive configuration studies.
@@ 4.5.1 Impact of PTQ Candidate Diversity
A central motivation of CALM is the premise that a single quantization method cannot optimally accommodate the distinct characteristics of every layer. To validate this, we systematically excluded individual quantization algorithms from the candidate pool and observed the resulting impact on model performance. As shown in table 3,experimental results demonstrate that retaining the full spectrum of candidate quantization methods yields superior performance across all evaluated models and tasks. The exclusion of any single algorithm invariably leads to a marked degradation in model fidelity.
@@
Table 3: Downstream task performance across five LLMs. Higher accuracy($\uparrow$) indicates better performance.(GSM=GSM8K, HE=HumanEval, HS=HellaSwag). ’w/o’ stands for ’without’.
For instance, on the Llama-3-8B model, our comprehensive strategy pool achieves a C4 PPL of 12.72. In contrast, removing SmoothQuant causes the C4 PPL to deteriorate significantly to 13.55, while excluding AWQ results in a drop in GSM8K reasoning accuracy from 74.33% to 72.85%.